To what pages must Narding open his book so that the product of the book's two facing pages is 5700? 17292?
Answer by tommyt3rd(5050)

The pages are consecutive so

n(n+1)=5700
n^2+n-5700=0
(n-76)(n+75)

page 75
The second problem is the same (has the same procedure for solution)
n(n+1)=17292
leads to page 131

:)